An owl is carrying a mouse to the chicks in its nest. It is 4.00 m west and 12.0 m above the center of the 30 cm diameter nest and is flying east at 3.00 m/s at an angle 27° below the horizontal when it accidentally drops the mouse. Will it fall into the nest? Find out by solving for the horizontal position of the mouse (measured from the point of release) when it has fallen the 12.0 m.             = m (from the point of release)
Answer #1

Apply kinematic equ tion along y direction

y - y0 = v0y t + 1/2 a t^2

0 = 12.0 -3.0*sin(27)*t - 0.5*9.81*t^2

solving qudratic equation

t=1.43 s

Apply ikinematic eqation along x direction

x = x0+ v0x t

= -4 +3.0*cos(27 )*1.43

=-0.178 m

radius of nest is r = d/2 = 0.30/2 = 0.15m

so the distance is = 0.178-0.15 = -0.028 m

the direction is west of the nest

No, the owl is not lucky. The mouse just misses the nest.